Group Selection Page HELP Controls Group ITDM Vacuum WP78 FXE SPB SQS SXP XO Migrated2Zulip Archived2GitlabPages Archived2PDF not2MigrateOrArchive MigratedByGroup
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip DET Sample Environment SPB_mig
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 649     Entry time: 17 Sep 2025, 09:40
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 17.09.2025 
General Announcement:
- F. Wolff-Fabris comments on the LIMP-25: for cw38 schedule sees: 
1. the PLC updates as main activity; 
2. Karabo upgrade in October with tests being carried out this week at MID; 
3. One pump in the KW20 circuit is being serviced from today (17.09) and thus no redundancy exists for one week. The KW20 water is supplied normally in this period.
- Two short power glitches were detected at DESY yesterday (17.09) morning. We are not aware of any disturbances at the campus but please let XO know if you were affected.

CTRL:
- Started MID migration last Monday and Tuesday. It was successful and no major issues happened, only few devices still missing; minor issues with Beckhoff. All identified and a fix to be done 
today/tomorrow.

DA:
- Nothing to report.

EEE:
- Nothing to report.

VAC:
- Not present.

XPD:
- Nothing to report.
- PLC was updated in SA1 and VC loop also. Few hours of pressure increase but still manageable and went fine.

DET:
- Nothing to report.

ITDM: 
- Nothing to report.

XRO:
- Not present.

LAS:
- Not present.

UNSYS:
- Everything is fine; Installation is ongoing.

SPB:
- Nothing to report.

FXE:
- Nothing to report.

MID:
- All fine at MID.

HED:
- Not present.

SCS:
- Not present.

SQS:
- Not present.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00